If 400 lb of 40% cream are mixed with 800 lb of milk, and if 300 lb of the mixture is mixed with 300 lb of 0% skim, and the final mixture tests 8% fat, what does the 800 lb of milk test?
Answer #1

Sot: 40016 of 40% cream – cream content - 40 x 400 = 160 16 milk Content = 60 x 400 - 240 lb Let the cream 7. in 850 eb of miwhen 300 lb of of skim & Added - cream content will be same as before in = (40+ 2x) eb Whereas, a milk content = 300 + (260-2